# DesCLIP: Robust Continual Adaptation via General Attribute Descriptions for Pretrained Vision-Language Models

### v1

@@ V Limitations
Observed in experimental trials, The effectiveness of our DesCLIP framework hinges on the language assistant (or real experts) being knowledgeable about the general attribute features of objects. This can pose challenges in certain applications that require indirect reasoning, such as StanfordCars [63], where it is difficult to accurately describe the representative features of a vehicle model associated with a specific license plate. Additionally, there are high standards for the quality of attribute descriptions; inappropriate prompts can introduce domain bias, ultimately hindering the ability of the Anchor-based embedding filter (AEF) to select highly relevant attribute features.
@@ VI Conclusions
Current research for VLM-based continual learning predominantly emphasizes to connect visual inputs and specific new-class text in downstream tasks, frequently neglecting the latent relationship between general knowledge and specialized knowledge for VLMs. In this paper, we propose DesCLIP, a framework that harnesses general attribute (GA) descriptions to enhance VLMs in establishing robust vision-GA-class text associations. By going beyond the traditional connections between visual inputs and class texts, DesCLIP employs a language assistant to generate candidates for attribute descriptions through tailored prompting. Additionally, we implemented an anchor-based embedding filter (AEF) to extract highly relevant description embeddings, which serve as paired text embeddings for instance matching (IM). Addtionally, we perform text embedding calibration (TEC) which allows for the progressive calibration of rudimentary text embeddings to align with representative GA representations. Our extensive experiments validate the effectiveness and advancements of DesCLIP, demonstrating its superior performance over existing pretrained and VLM-based continual learning methods.

### v2

@@ V-C Limitations
Observed in experimental trials, the effectiveness of our DesCLIP hinges on the language assistant (or real experts) being knowledgeable about the general attribute features of objects. This can pose challenges in certain applications that require indirect reasoning, where it is difficult to accurately describe the representative features of a vehicle model associated with a specific license plate. Furthermore, our method may face limitations when expert knowledge is unavailable or when encountering OOD visual objects that are inherently difficult to describe. In such scenarios, approaches that mine patterns solely from single-modal visual representations (e.g., GDA) could be more effective, as they bypass the dependency on precise attribute descriptions and rely instead on direct visual regularities.
@@ VI Conclusions
Current research on VLM-based continual learning primarily focuses on acquiring downstream knowledge through opaque parameter updates and implicit model adjustments, often overlooking the structured relationship between general and task-specific knowledge. In this paper, we propose DesCLIP, a framework that harnesses general attribute (GA) descriptions to enhance VLMs in establishing robust vision-GA-class text associations. By going beyond the traditional connections between visual inputs and class texts, DesCLIP employs a language assistant to generate candidates for attribute descriptions through tailored prompting. Additionally, we implement an anchor-based embedding filter (AEF) to extract highly relevant description embeddings, which serve as paired text embeddings for instance matching (IM). Addtionally, we perform text embedding calibration (TEC) which allows for the progressive calibration of rudimentary text embeddings to align with representative GA representations. Our extensive experiments validate the effectiveness and advancements of DesCLIP, demonstrating its superior performance over existing pretrained and VLM-based continual learning methods.